Circular No. 8447

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004gm IN MCG –02-33-80 Further to IAUC 8445, J. Burket and W. Li report the LOSS discovery

of an apparent supernova (mag 15.9) on an unfiltered KAIT image taken on Nov. 30.55 UT. The new object, designated SN 2004gm, is located at α = 12^{h}59^{m}40^{s}.53, δ = −14^{o}58′02′′.4 (equinox 2000.0), which is 9′′.8 east and 0′′.4 south of the center of MCG –02-33-80; nothing was visible at this position on an image taken on June 5.21 (limiting mag 19.5).

T. Matheson, National Optical Astronomy Observatory; and P. Challis, A. Garg, M. Modjaz, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 420–970 nm) of SN 2004gm, obtained by Garg and Kirshner on Dec. 1.34 UT with the Magellan Baade 6.5-m telescope (+ IMACS spectrograph), shows it to be a type-Ia supernova, several weeks past maximum.

SUPERNOVAE 2004gg, 2004gh, 2004gi G. Folatelli, M. Hamuy, N. Morrell, and M. Phillips, Carnegie Super-

nova Project, report that CCD spectra (range 360–900 nm) of SN 2004gg, SN 2004gh, and SN 2004gi (IAUC 8444) were obtained by M. Roth on Nov. 25 and 26 UT, using the Magellan II (Clay) telescope (+ LDSS-2 spectrograph) at Las Campanas Observatory. The spectrum of 2004gg closely resembles that of the peculiar type-II SN 1987K, ≈ 1 week after maximum light. SN 2004gh is a type-II supernova, ∼ 1 month after explosion, showing a prominent Hα P-Cyg profile and an expansion velocity of 4000 km/s, as inferred from the absorption minimum of Hβ and assuming the recession velocity of the host galaxy to be 3662 km/s, as given in the NED database. The spectrum of SN 2004gi is that of a type-Ia supernova, ∼ 25 days after maximum light.

V574 PUPPIS N. M. Ashok and D. P. K. Banerjee, Physical Research Laboratory,

Ahmedabad, report that near-infrared JHK spectroscopy (range 1.08–2.35 μm) of V574 Pup, taken on Nov. 26.98 UT with the Mt. Abu 1.2-m telescope (+ PRL Near-Infrared Imaging Spectrometer), shows strong H I emission lines of Paβ, Paγ, Brγ, and Brackett 11 to 17. Prominent O I lines at 1.1288 and 1.3166 μm are seen, as are a blend of N I and C I lines in the region around 1.175 and 1.25 μm.

2004 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green